# The Tension of Space as Dark Energy: A No Geometric Sequestering Theorem for Minimal Sequesters

### v2

@@ 2.3 Space as the Three Brane
It is important to mention here that we do not assume that physical space is embedded in a higher dimensional ambient manifold. The elastic membrane language in this paper is an intrinsic description and all observables are built from the intrinsic metric and its derived tensors. Introducing an embedding is possible as a change of variables but it adds redundant structure that is not needed for the dynamics we study. An embedding also introduces an arbitrary ambient metric that does not enter any measurement performed by observers confined to the physical spacetime. Our equations of motion are fully determined by intrinsic geometry, so the theory remains sensible without any reference to extrinsic data. For these reasons we avoid committing to an embedding and keep the framework intrinsic from the outset.
3 A No-Geometric-Neutralization Theorem for Sequestering Four–Forms
We now work in natural units $\hbar=c=1$ in $3+1$ dimensions and introduce the sequester action which includes the usual Einstein–Hilbert term, a fundamental $3$-form Nambu-Goto action for space (essentially treating it as an elastic membrane), a matter sector that couples to a globally scaled metric through a single global variable $\lambda$ and two four-form sectors $F_{4}$ and $\hat{F}_{4}$ that enforce global constraints, see [61, 62, 63, 64, 65] for the detailed framework. The action has the following form,
$S=$ $\int d^{4}x\,\sqrt{-g}\,\Big(\frac{M_{*}^{2}}{2}\,R-T_{s}-\Lambda\Big)+\int d^{4}x\,\sqrt{-g}\,L_{m}(\lambda^{-2}\,g_{\mu\nu},\psi)$
$+\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)\,\int F_{4}+\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\int\hat{F}_{4}$ (3.1)
with $M_{*}^{2}=1/8\pi G$. Note that we are working in a convention where the units of $\Lambda$ are $\Lambda]=\rm mass^{4}$ (energy density). The matter fields are collected into $\psi$ and the scale $\mu$ is a fixed reference scale. A note on notation, content of the action and minimal assumptions are in order,
a) We work with a Lorentzian metric $g_{\mu\nu}$ of signature (+++)(-+++). The Levi–Civita connection is metric compatible and curvature tensors have the usual definitions.
b) The total action includes the following pieces:
$\bullet$ Gravity: $\int\sqrt{-g}\,\frac{M_{*}^{2}}{2}\,R$.
$\bullet$ Geometric tension of space: $-\,T_{s}\int\sqrt{-g}$. One may think of $T_{s}$ as a constant that multiplies the physical spacetime volume. It is not a field and has no kinetic term and it only encodes an NG-like tension of the underlying elastic space. It is purely geometric in the sense that it couples only through $\sqrt{-g}$.
$\bullet$ A bare cosmological variable $\Lambda$ that appears as $-\Lambda\int\sqrt{-g}$.
$\bullet$ Matter with a global Weyl scaling through $\lambda$:
$S_{m}[g,\psi;\lambda]\equiv\int\sqrt{-g}\,L_{m}(\lambda^{-2}\,g_{\mu\nu},\psi)$ (3.2)
The factor $\lambda^{-2}g_{\mu\nu}$ ensures that the matter sector sees a metric that is globally Weyl related to $g_{\mu\nu}$. The variable $\lambda$ is not a field with spacetime dependence rather it is a single global variable. We will see this from the four-form sector later.
$\bullet$ Two top-form sectors,
$S_{4\text{-forms}}=\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)\int F_{4}+\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\int\hat{F}_{4}$ (3.3)
where $F_{4}=dA_{3}$ and $\hat{F}_{4}=d\hat{A}_{3}$ with $A$ and $\hat{A}$ being three form potentials (see [54, 55, 57, 58] for a working understanding of three forms). The integrals $\int F_{4}$ and $\int\hat{F}_{4}$ do not use the volume factor $\sqrt{-g}$. In essence, they are purely topological and their values depend on global fluxes and boundary conditions, not on the local metric.
c) Assumptions about the functions $\sigma$ and $\hat{\sigma}$: We assume that they are smooth and monotone enough so that the equations we derive can be inverted to view $\Lambda$ and $\lambda$ as global constants fixed by fluxes. We do not need stronger properties at this stage. We also implicitly assume boundary conditions that render the variational problem well posed for $A_{3}$ and $\hat{A}_{3}$.
d) Variation conventions: Variation with respect to $g^{\mu\nu}$ defines the total stress tensor as usual. Variation with respect to $\psi$ gives matter equations of motion that we do not need explicitly. Variation with respect to $A_{3}$ and $\hat{A}_{3}$ gives four-form equations that enforce global constraints. Variation with respect to $\Lambda$ and $\lambda$ gives global relations because the four-forms enter as Lagrange multipliers as we will demonstrate below.
In all, this is the standard sequestering setup except there is an additional NG-like term for space.
We now proceed to show that under a constant shift of the matter Lagrangian $L_{m}$ by $\Delta V$ the combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ is invariant point-wise . The geometric tension $T_{s}$ is untouched by the four-form constraints because it does not couple to $\lambda$ and does not appear in the matter trace. Therefore the effective cosmological term left in the local Einstein equations equals $T_{s}$. A four-form sequestering mechanism of this type cannot neutralize the NG-like geometric tension of space.
Proof:
1.) Varying the action with respect to the three-form potentials:
Consider the term $\sigma(\Lambda/\mu^{4})\int F_{4}$, where $F_{4}=dA_{3}$. Under a variation of $A_{3}$ with compact support we have,
$\delta\left(\int\sigma\,F_{4}\right)=\int\sigma\,\delta F_{4}=\int\sigma\,d(\delta A_{3})=\int d\big(\sigma\,\delta A_{3}\big)-\int d\sigma\wedge\delta A_{3}$ (3.4)
The boundary term is $\int d(\sigma\,\delta A_{3})$. We set it to zero by appropriate boundary conditions. The remaining term is $\int d\sigma\wedge\delta A_{3}$ and given that $\delta A_{3}$ is arbitrary in the interior, the Euler–Lagrange equation is,
$d\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)=0$ (3.5)
Now because $\sigma$ depends on spacetime only through $\Lambda$, this implies that $d(\Lambda/\mu^{4})=0$ provided $\sigma^{\prime}\neq 0$. Therefore $\Lambda$ is constant in spacetime.
Repeating the same steps for $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$, we have after varying with respect to $\hat{A}_{3}$,
$d\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)=0$ (3.6)
Assuming $\hat{\sigma}^{\prime}\neq 0$ we conclude that $d(\lambda^{4}/\mu^{4})=0$. Hence $\lambda$ is constant in spacetime and since $\lambda>0$ we can take $\lambda$ itself to be a global constant.
Therefore, the four-form sectors force $\Lambda$ and $\lambda$ to be global constants. That is, they do not propagate and they are numbers fixed by boundary data and fluxes.
2.) Varying the action with respect to the metric and identifying the local Einstein equations:
We now compute the variation of each metric dependent term with respect to $g^{\mu\nu}$ where we will be using the identity $\delta\sqrt{-g}=-\dfrac{1}{2}\sqrt{-g}\,g_{\mu\nu}\,\delta g^{\mu\nu}$. The variation of the Einstein–Hilbert part gives the usual $M_{*}^{2}G_{\mu\nu}$ on the left hand side of the field equations.
The variation of $T_{s}\int\sqrt{-g}$ gives,
$\delta\big(-T_{s}\int\sqrt{-g}\big)=-T_{s}\int\delta\sqrt{-g}=\frac{T_{s}}{2}\int\sqrt{-g}\,g_{\mu\nu}\,\delta g^{\mu\nu}$ (3.7)
This means the geometric tension contributes a stress tensor equal to $T_{s}g_{\mu\nu}$ with a minus sign when moved to the right hand side of Einstein equations.
The variation of $\Lambda\int\sqrt{-g}$ is identical in structure and it gives a negative contribution $\Lambda g_{\mu\nu}$ to the right hand side of the field equations as well.
Now we vary the matter action with respect to the metric, which reads,
$S_{m}[g,\psi;\lambda]=\int\sqrt{-g}\,\,L_{m}(\lambda^{-2}\,g,\psi)$ (3.8)
We define the physical matter stress tensor by the standard definition,
$T^{(m)}_{\mu\nu}\equiv-\frac{2}{\sqrt{-g}}\,\frac{\delta S_{m}}{\delta g^{\mu\nu}}$ (3.9)
The above definition ensures that the matter equations are consistent with diffeomorphism invariance and with the Bianchi identities. The full variation of the matter part yields,
$\delta S_{m}=-\frac{1}{2}\int\sqrt{-g}\,T^{(m)}_{\mu\nu}\,\delta g^{\mu\nu}$ (3.10)
Collecting all terms, the total variation with respect to $g^{\mu\nu}$ gives,
$\delta S=\int\sqrt{-g}\,\Big[\frac{M_{*}^{2}}{2}\,G_{\mu\nu}+\frac{1}{2}\big(T_{s}+\Lambda\big)g_{\mu\nu}-\frac{1}{2}\,T^{(m)}_{\mu\nu}\Big]\,\delta g^{\mu\nu}$ (3.11)
Stationarity for arbitrary $\delta g^{\mu\nu}$ gives,
$\frac{M_{*}^{2}}{2}\,G_{\mu\nu}+\frac{1}{2}\big(T_{s}+\Lambda\big)g_{\mu\nu}-\frac{1}{2}\,T^{(m)}_{\mu\nu}=0$ (3.12)
Multiplying by 2, the local Einstein field equations read,
$M_{*}^{2}\,G_{\mu\nu}=T^{(m)}_{\mu\nu}-\big(T_{s}+\Lambda\big)\,g_{\mu\nu}$ (E)
It is important to note that the four-form terms do not contribute to the metric equations. They are metric independent in the form we wrote, since $\int F_{4}$ and $\int\hat{F}_{4}$ do not carry $\sqrt{-g}$ or $g^{\mu\nu}$. Hence they add no local stress tensor.
So, we now have the local field equations where we see that the geometric tension $T_{s}$ and the variable $\Lambda$ both appear as volume terms and therefore enter as $(\Lambda+T_{s})\,g_{\mu\nu}$. The matter sector contributes its usual stress tensor $T^{(m)}_{\mu\nu}$.
3.) Varying the action with respect to $\Lambda$:
Only two parts of the action (3.1) depend on $\Lambda$. The volume term $\Lambda\int\sqrt{-g}$ and the top-form term $\sigma(\Lambda/\mu^{4})\int F_{4}$. Therefore, computing the variation and treating $\Lambda$ as a global number, we find,
$\delta S\big|_{\Lambda}=-\delta\Lambda\int d^{4}x\,\sqrt{-g}+\sigma^{\prime}\Big(\frac{\Lambda}{\mu^{4}}\Big)\,\frac{\delta\Lambda}{\mu^{4}}\,\int F_{4}$ (3.13)
Setting the equation above this to zero for arbitrary $\delta\Lambda$, we obtain the first global constraint,
$\int d^{4}x\,\sqrt{-g}=\frac{\sigma^{\prime}\left(\Lambda/\mu^{4}\right)}{\mu^{4}}\,\int F_{4}$ (C1)
This equation tells us that the spacetime four-volume is proportional to the four-form flux weighted by $\sigma^{\prime}$. Note that it does not fix the local metric rather it only relates global numbers.
With this, we have one global relation that involves $\Lambda$ but no local fields directly.
4.) Varying the action with respect to $\lambda$:
Again, only two parts of the action (3.1) depend on $\lambda$, which are the matter action $S_{m}$ and the top-form term $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$. First consider the matter action which we compute $\delta S_{m}$ under a small variation of the global $\lambda$.
$\delta S_{m}\bigg|_{\lambda}=\int\sqrt{-g}\,\frac{\partial L_{m}(\lambda^{-2}\,g,\,\psi)}{\partial(\lambda^{-2}g_{\mu\nu})}\,\delta(\lambda^{-2}g_{\mu\nu})$ (3.14)
Let $h_{\mu\nu}\equiv\lambda^{-2}g_{\mu\nu}$ and its variation under a variation of $\lambda$ at fixed $g$ is,
$\delta h_{\mu\nu}=\delta(\lambda^{-2})\,g_{\mu\nu}=-2\,\lambda^{-3}\,\delta\lambda\,g_{\mu\nu}$ (3.15)
So we have,
$\int\sqrt{-g}\;\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\;\delta(h_{\mu\nu})=-2\,\int\sqrt{-g}\;\lambda^{-3}\,\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}\,\delta\lambda$ (3.16)
In order to obtain $\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}$ , we first vary $S_{m}$ with respect to the metric tensor $g_{\mu\nu}$, which gives,
$\delta S_{m}=-\int\sqrt{-g}\;\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\alpha\mu}\,g_{\beta\nu}\delta g^{\mu\nu}$ (3.17)
and use the definition of the matter stress energy tensor (3.9), which gives,
$T_{\mu\nu}^{(m)}=2\,\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\alpha\mu}\,g_{\beta\nu}$ (3.18)
Taking the trace, we find,
$T^{(m)}=g^{\mu\nu}\,T_{\mu\nu}^{(m)}=2\,\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\mu\nu}$ (3.19)
Solving for $\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}$, we obtain,
$\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}g_{\mu\nu}=\frac{\lambda^{2}\,T^{(m)}}{2}$ (3.20)
Substituting into equation (3.16), we get,
$\int\sqrt{-g}\;\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\;\delta(h_{\mu\nu})=-\frac{\delta\lambda}{\lambda}\int\sqrt{-g}\,T^{(m)}$ (3.21)
Using (3.21) in (3.14), we obtain,
$\delta S_{m}\bigg|_{\lambda}=-\frac{\delta\lambda}{\lambda}\int\sqrt{-g}\,T^{(m)}$ (3.22)
Now we vary the top-form term $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$, which gives,
$\delta\Big[\hat{\sigma}\!\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\int\hat{F}_{4}\Big]=\hat{\sigma}^{\prime}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\frac{4\lambda^{3}\,\delta\lambda}{\mu^{4}}\,\int\hat{F}_{4}$ (3.23)
Writing both pieces together, that is equations (3.21) and (3.23), the full variation of $S_{m}$ with respect to the global $\lambda$ then reads,
$\frac{\delta\lambda}{\lambda}\,\int\sqrt{-g}\,T^{(m)}+\hat{\sigma}^{\prime}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\frac{4\lambda^{3}\,\delta\lambda}{\mu^{4}}\,\int\hat{F}_{4}=0$ (3.24)
Canceling $\delta\lambda$ which is arbitrary but not zero and multiplying by $\lambda$, we obtain the second global constraint,
$\int d^{4}x\,\sqrt{-g}\,T^{(m)}=\frac{4\,\lambda^{4}}{\mu^{4}}\,\hat{\sigma}^{\prime}\left(\frac{\lambda^{4}}{\mu^{4}}\right)\,\int\hat{F}_{4}$ (C2)
We now have a second global relation that connects the spacetime integral of the matter trace to the flux of $\hat{F}_{4}$. The geometric tension $T_{s}$ does not appear because it does not depend on $\lambda$.
5.) Eliminating fluxes to relate $\Lambda$ to the spacetime average of $T^{(m)}$:
Equations (C1) and (C2) contain the fluxes $\int F_{4}$ and $\int\hat{F}_{4}$. These fluxes are constant numbers fixed by boundary conditions. We define the spacetime average of a quantity $X$ as
$\langle X\rangle=\dfrac{\int\sqrt{-g}\,X}{\int\sqrt{-g}}$ (3.25)
Combining (C1) and (C2) we may then write,
$\langle T^{(m)}\rangle=4\,\lambda^{4}\,\frac{\hat{\sigma}^{\prime}(\lambda^{4}/\mu^{4})}{\sigma^{\prime}(\Lambda/\mu^{4})}\,\frac{\int\hat{F}_{4}}{\int F_{4}}$ (3.26)
[…]

### v3

@@ 2.3 Space as a $D_{3}$ Brane
Furthermore, it is important to mention here that we do not assume that physical space is embedded in a higher dimensional ambient manifold. The elastic membrane language in this paper is an intrinsic description and all observables are built from the intrinsic metric and its derived tensors. Introducing an embedding is possible as a change of variables but it adds redundant structure that is not needed for the dynamics we study. An embedding also introduces an arbitrary ambient metric that does not enter any measurement performed by observers confined to the physical spacetime. Our equations of motion are fully determined by intrinsic geometry, so the theory remains sensible without any reference to extrinsic data. For these reasons we avoid committing to an embedding and keep the framework intrinsic from the outset.
The construction developed so far captures only the uniform tension sector of the intrinsic membrane. However, once the membrane of space is allowed to support an internal hidden $U(1)$ gauge field, the Nambu–Goto description is no longer the most complete local nonlinear effective action. The natural generalization is then of Dirac–Born–Infeld type. In the limit that the hidden gauge field is switched off, the DBI action reduces to the Nambu–Goto volume term. When the gauge field is present, it resums the nonlinear response of the membrane to finite worldvolume fluxes and provides a finite-field completion of the Maxwell regime. We therefore turn next to the DBI completion of the intrinsic membrane picture and ask what kind of hidden magnetic sectors it admits.
2.4 The DBI Completion of the Elastic Space Picture
A convenient starting point is the four-dimensional DBI gauge action (we work in units where $\hbar=c=1$),
$S_{\rm DBI}=-T_{s}\int d^{4}x\,\sqrt{-\det\left(g_{\mu\nu}+\lambda_{h}F_{\mu\nu}\right)}\,.$ (2.31)
Factoring out the metric determinant gives,
$S_{\rm DBI}=-T_{s}\int d^{4}x\,\sqrt{-g}\,\sqrt{\det\left(\delta^{\mu}{}_{\nu}+\lambda_{h}F^{\mu}{}_{\nu}\right)}\,.$ (2.32)
Since $F_{\mu\nu}$ is antisymmetric, the determinant identity in four dimensions yields,
$\det\left(\delta^{\mu}{}_{\nu}+\lambda_{h}F^{\mu}{}_{\nu}\right)=1+\frac{\lambda_{h}^{2}}{2}F_{\mu\nu}\,F^{\mu\nu}-\frac{\lambda_{h}^{4}}{16}(F_{\mu\nu}\tilde{F}^{\mu\nu})^{2}\,.$ (2.33)
Therefore,
$S_{\rm DBI}=-T_{s}\int d^{4}x\,\sqrt{-g}\,\sqrt{1+\frac{\lambda_{h}^{2}}{2}F_{\mu\nu}\,F^{\mu\nu}-\frac{\lambda_{h}^{4}}{16}(F_{\mu\nu}\tilde{F}^{\mu\nu})^{2}}\,.$ (2.34)
Writing $\lambda_{h}=\beta_{h}^{-1}$, one obtains,
$S_{\rm DBI}=-T_{s}\int d^{4}x\,\sqrt{-g}\,\sqrt{1+\frac{1}{2\beta_{h}^{2}}\,F_{\mu\nu}\,F^{\mu\nu}-\frac{1}{16\beta_{h}^{4}}\,(F_{\mu\nu}\tilde{F}^{\mu\nu})^{2}}\,.$ (2.35)
Note that the equation above in weak fields reduces to,
$S_{\rm DBI}=-T_{s}\int d^{4}x\,\sqrt{-g}\,\left[1+\frac{1}{4\,\beta_{h}^{2}}\,F_{\mu\nu}\,F^{\mu\nu}\right]\,.$ (2.36)
When $\beta_{h}^{2}=T_{s}$ as in the Maxwell normalization, we have,
$S_{\rm DBI}=-\int d^{4}x\,\sqrt{-g}\,\left[T_{s}+\frac{1}{4}\,F_{\mu\nu}\,F^{\mu\nu}\right]\,$, (2.37)
with the Lagrangian density,
$\mathcal{L}_{\rm DBI}=-T_{s}-\frac{1}{4}F_{\mu\nu}\,F^{\mu\nu}\,.$ (2.38)
At this point, the minimal physical content of the construction becomes clear. Once space is treated as an effective DBI membrane, the theory already contains two basic ingredients. The first is the vacuum-like or cosmological contribution $T_{s}$, which in the present framework is identified with the residual observed vacuum sector. The second is an abelian gauge sector arising from the DBI structure itself, whose weak-field limit is the $U(1)$ field appearing above. In this work, we regard that $U(1)$ as hidden. The reason is simple. Nothing in the bare effective action requires it to be identified with ordinary electromagnetism, and in particular the visible matter content and coupling assignments do not follow automatically from the minimal DBI sector alone. One may certainly imagine more elaborate ultraviolet completions in which Standard Model gauge fields and matter are engineered from intersecting or stacked brane configurations, but such constructions are highly model dependent and often rather contrived. Our purpose here is more modest. We only use the minimal DBI completion as an intrinsic effective description of space, which naturally furnishes a vacuum term together with a hidden $U(1)$ sector, without assuming that this sector is the electromagnetic field of the Standard Model.
3 Late-Time Running of the Tension of Space
We now consider a late-time microscopic origin for the running of the space tension. The purpose of the present section is not to provide a complete ultraviolet model of hidden-sector interaction. Rather, we construct a proof of concept effective description in which the intrinsic hidden gauge sector already suggested by the DBI completion of space develops, at late times, a defect-like reservoir that can exchange energy with the vacuum-like tension sector. We consider here an additional hidden-sector transition at late times, whose detailed microphysics we do not attempt to derive here, but whose coarse-grained cosmological effects can still be studied consistently.
3.1 Hidden $U(1)\to\mathbb{Z}_{n}$ breaking and Flux-Tube Defects
We now extend the intrinsic hidden $U(1)$ gauge sector of space, already introduced in the DBI description, by a complex scalar field $\Phi$ carrying charge $n$ under that same hidden Abelian symmetry. At the level of a minimal effective description, the hidden sector is modeled by extending the DBI Lagrangian of space, equation (2.38), so as to include a hidden Higgs field. The corresponding effective Lagrangian is
$\mathcal{L}_{\rm hid}=-\frac{1}{4}Z(\sigma)\,F_{\mu\nu}\,F^{\mu\nu}+|D_{\mu}\Phi|^{2}-\frac{\lambda_{\Phi}}{4}\left(|\Phi|^{2}-v_{\Phi}^{2}(\sigma)\right)^{2}\,$, (3.1)
with,
$D_{\mu}\Phi=\left(\partial_{\mu}-i\,ne_{h}A_{\mu}\right)\Phi\,.$ (3.2)
Here $A_{\mu}$ is the hidden gauge field, $e_{h}$ is the hidden gauge coupling, and $\sigma$ is a slow collective variable that will be used below to parameterize the dynamical part of the space-tension sector. The functions $Z(\sigma)$ and $v_{\Phi}(\sigma)$ allow the hidden sector to communicate with the tension sector at the effective level.
When the hidden Higgs field condenses, at a late enough time in cosmic history (as discussed in Sec. 4.2),
$\langle\Phi\rangle=\frac{v_{\Phi}}{\sqrt{2}}\neq 0\,$, (3.3)
the continuous hidden gauge symmetry is not necessarily destroyed completely. Because $\Phi$ carries charge $n$, the vacuum remains invariant under the discrete subgroup $\mathbb{Z}_{n}\subset U(1)_{h}$ [60, 61]. At the effective level, the late-time symmetry-breaking pattern is therefore,
$U(1)_{h}\longrightarrow\mathbb{Z}_{n}\,.$ (3.4)
Owing to this condensation, the late-time hidden sector enters a topologically nontrivial broken phase in which localized flux-tube defects can form, see [62, 63, 64, 65]. The basic mechanism is standard. During the transition, different causal regions need not select the same phase of the hidden complex scalar $\Phi$. As a result, the phase of $\Phi$ can acquire a nontrivial winding around certain closed loops in space. Such a winding cannot be removed continuously while remaining everywhere in the broken phase. At least along some line-like locus, the scalar magnitude must therefore be driven back toward zero so that the phase becomes ill-defined there. These line-like loci are precisely the cores of the hidden flux tubes.
The gauge field then adjusts so as to maintain finite energy away from the core. In particular, the covariant gradient energy of the Higgs field requires the phase winding of $\Phi$ to be compensated by an azimuthal gauge configuration, so that the hidden magnetic flux becomes trapped inside the defect core rather than remaining as a freely propagating long-range field. Equivalently, once the hidden $U(1)$ sector is Higgsed to a residual discrete subgroup, the gauge boson acquires a mass in the broken phase, and the bulk expels magnetic flux in the usual Meissner-like manner [62, 66, 67]. The flux is therefore energetically forced into narrow, string-like tubes. For a winding number $k$, the trapped flux is quantized, schematically as,
$\Phi_{B}=\oint A_{\theta}\,r\,d\theta=\frac{2\pi k}{ne_{h}}\,,\qquad k\in\mathbb{Z}\,$, (3.5)
so that the resulting objects are stable localized carriers of hidden magnetic flux.
In this way, the late-time symmetry breaking does not merely leave behind a bath of free hidden magnetic radiation. Rather, it reorganizes the relevant hidden energy into an extended network of flux-tube defects whose microscopic width is set by the inverse hidden Higgs and gauge masses, while their lengths may be macroscopic on cosmological scales. The detailed nonequilibrium dynamics of the transition, including defect formation, intercommutation, loop production, and the subsequent network evolution [64, 65], are not required at the phenomenological level. For the present proof-of-concept analysis, it is sufficient to treat the magnetic flux network in coarse-grained form as a hidden flux-tube reservoir with an effective energy density $\rho_{\rm flux}$ and an effective equation-of-state parameter $w_{\rm flux}$.
Once such a reservoir is present, the next question is how it communicates with the vacuum-like sector associated with the intrinsic tension of space. In particular, if the hidden defect network can exchange energy with the dark-energy sector, then the total tension can no longer remain purely rigid at late times. This motivates a decomposition of the space tension into a constant background piece and a dynamical contribution, from which a time-dependent effective equation of state for dark energy naturally emerges.
3.2 Splitting the Tension Sector and the Effective Equation of State of Dark Energy
If we assume that the energy carried by the magnetic flux tube defects produced due to the breaking of $U(1)$ gauge symmetry by the hidden Higgs field can be exchanged with the vacuum sector, then the spatial tension $T_{s}$ identified in Sec. 2 splits into a constant background piece and a dynamical transferred piece,
$T(t)=T_{s}+\delta T_{s}(t)\,,\qquad\dot{T}_{s}=0\,.$ (3.6)
Here $T_{s}$ denotes as before the rigid background contribution that survives after the ultraviolet physics has played its role. It is identified with the dominant constant part of the observed dark-energy scale as mentioned earlier. By contrast, $\delta T_{s}$ represents the late-time correction induced by energy exchange with the hidden flux-tube sector. Only $\delta T_{s}$ is allowed to participate in the exchange. This is physically important as the constant background $T_{s}$ is the baseline vacuum sector itself and is understood as a built-in background parameter inherited from hard ultraviolet complete physics. In that sense it is rigid and cannot be treated as a reservoir that either drains energy into other sectors or receives energy back from them. It is simply the irreducible vacuum baseline that remains after the ultraviolet problem has already been resolved.
What can evolve is only an additional contribution deposited on top of this baseline by later hidden-sector dynamics. Therefore $\delta T_{s}$ denotes the excess energy temporarily stored in the tension sector. Any subsequent relaxation, unloading, or transfer should then be understood as acting only on $\delta T_{s}$, not on the underlying vacuum piece $T_{s}$ itself. The late-time dynamics does not remove the intrinsic vacuum tension of space. It only perturbs the total tension away from its rigid baseline and may later allow that excess contribution to decay.
At the background level, we continue to identify the dark-energy density with the total space tension,
$\rho_{\rm DE}(t)\;\equiv\;T(t)\;=\;T_{s}+\delta T_{s}(t)\,.$ (3.7)
Since $\rho_{\rm DE}$ changes with time, or equivalently the scale factor $a$, its effective equation of state is also time dependent and reads (see Appendix A),
$w_{\rm eff}(a)=-1-\frac{1}{3}\frac{d\ln T}{d\ln a}=-1-\frac{1}{3\bigl(T_{s}+\delta T_{s}\bigr)}\frac{d\delta T_{s}}{d\ln a}\,.$ (3.8)
One can therefore see that if $\delta T_{s}$ changes with the scale factor, then $w_{\rm eff}$ changes as well, giving a dynamical equation of state for dark energy.
Equation (3.8) makes the central point of the present framework explicit. Any late-time exchange that modifies $\delta T_{s}$ immediately induces a departure of $w_{\rm eff}$ from 1-1, even though the rigid background contribution $T_{s}$ itself remains untouched. In this sense, the running of dark energy is not attributed to the erosion of the vacuum baseline, but to the evolution of an additional tension component deposited on top of it by hidden-sector dynamics. What remains is to represent this exchange in a simple local effective description. This is the purpose of the next subsection.
3.3 Pedagogical Lagrangian for the coupled Tension and Flux Sectors
The splitting introduced above is sufficient to define the background dark-energy sector and its effective equation of state, but it does not yet specify a local field-theoretic channel through which the hidden flux reservoir communicates with the dynamical part of the tension. To make that interaction more concrete, it is useful to introduce a simple pedagogical effective action in which the transferred part of the tension is encoded by a slowly varying scalar degree of freedom [35, 68]. This provides a compact local framework for organizing the couplings between the tension sector, the hidden gauge field, and the hidden Higgs field.
To write a local effective action for a dynamical exchange, it is convenient to represent the transferred part of the tension sector by a slowly varying scalar degree of freedom $\sigma$ such that,
$T(\sigma)=T_{s}+\delta T_{s}(a,\sigma)\,.$ (3.9)
A simple effective action follows from augmenting equation (3.1) with a term to encode the dynamics of the $\sigma$ sector and replace $T_{s}$ with $T=T_{s}+\delta T_{s}$,
$S=\int d^{4}x\,\sqrt{-g}\,\Bigg[-\Bigl(T_{s}$ $+\delta T_{s}(a,\sigma)\Bigr)-\frac{1}{4}Z(\sigma)\,F_{\mu\nu}\,F^{\mu\nu}$ (3.10)
$+|D_{\mu}\Phi|^{2}-\frac{\lambda_{\Phi}}{4}\left(|\Phi|^{2}-v_{\Phi}^{2}(\sigma)\right)^{2}-\frac{K_{\sigma}}{2}\,(\partial^{\mu}\sigma\,\partial_{\mu}\sigma)\Bigg]\,$,
where the quantity $K_{\sigma}$ is a constant. Furthermore, $\delta T_{s}(\sigma)$ makes explicit that only the dynamical correction to the tension is carried by $\sigma$, while $T_{s}$ remains constant. The couplings $Z(\sigma)$ and $v_{\Phi}(\sigma)$ provide an effective communication channel between the hidden flux-tube sector and the tension sector.
As mentioned earlier, the present treatment is intentionally phenomenological. We do not attempt to solve the full defect-formation and string-network dynamics from the field theory above. Instead, we move to a coarse-grained background description.
3.4 Coarse-grained Continuity Equations
Let $\rho_{\rm flux}$ denote the coarse-grained energy density of the hidden flux-tube network, and let its effective pressure be parameterized by,
$p_{\rm flux}=w_{\rm flux}\,\rho_{\rm flux}\,.$ (3.11)
For an almost frozen string-like reservoir one expects $w_{\rm flux}$ to be negative, while for a more rapidly evolving network it may be closer to zero. In the present proof-of-concept treatment we leave $w_{\rm flux}$ as an effective parameter.
We now assume that the transferred part of the tension sector and the flux-tube reservoir exchange energy through a bidirectional channel. The corresponding continuity equations are taken to be (see Appendix B),
$\dot{\rho}_{\rm flux}+3\,H\bigl(1+w_{\rm flux}\bigr)\rho_{\rm flux}$ $=-\,Q\,$, (3.12)
$\dot{\delta T_{s}}$ $=Q\,.$ (3.13)
where $Q$ is the amount of energy density exchanged per unit time. The exchange term is modeled phenomenologically as,
$Q=\Gamma H\left(\rho_{\rm flux}-\kappa\,T_{s}\right)\,,\qquad\Gamma\,>\,0\,,\qquad 0\,\leq\,\kappa\,\leq\,1\,.$ (3.14)
where $\Gamma$ is a dimensionless coupling parameter. This form captures the simplest balanced communication channel between the two dynamical sectors. It should be mentioned here that the phenomenological exchange law introduced above was written in its minimal form by taking the coupling parameter $\Gamma$ to be constant. However, once the transferred part of the tension sector is encoded by the scalar degree of freedom $\sigma$ as discussed in Sec. 3.3, it is natural to expect that the efficiency of the exchange channel itself may depend on the state of that sector. This expectation is reinforced by the fact that the same scalar already controls the couplings $Z(\sigma)$ and $v_{\Phi}(\sigma)$ in the local effective action. However, here we keep the analysis simple and treat $\Gamma$ as a constant. When the hidden flux-tube reservoir dominates over the stored dynamical correction to the space tension, that is when
$\rho_{\rm flux}\,>\,\kappa\,T_{s}\,$, (3.15)
one has Q> 0Q\,>\,0, so energy flows from the hidden defect sector into the space tension. By contrast, when
$\rho_{\rm flux}\,<\,\kappa\,T_{s}\,$, (3.16)
one has Q< 0Q\,<\,0, so the excess dynamical part of the tension sector relaxes back into the hidden flux sector. Importantly, the constant background piece $T_{s}$ never participates in this exchange.
Substituting (3.14) into (3.12) and (3.13), the coupled background system becomes
$\dot{\rho}_{\rm flux}+3H\bigl(1+w_{\rm flux}\bigr)\rho_{\rm flux}$ $=-\,\Gamma H\left(\rho_{\rm flux}-\kappa\,T_{s}\right)\,$, (3.17)
$\dot{\delta T_{s}}$ $=\Gamma H\left(\rho_{\rm flux}-\kappa\,T_{s}\right)\,.$ (3.18)
If we further assume that the energy density of the flux-tube-network is some fraction $\eta$ of the residual vacuum energy $T_{s}$, then,
$\rho_{\rm flux}=\eta(t)\,T_{s}\,,\qquad 0\,\leq\,\eta(t)\,\leq\,1$ (3.19)
which gives,
[…]